Ionizing Radiation

1.      Radon 222 is released from the earth and can become trapped in buildings. Suppose the concentration measured in a college classroom building is 1500 pCi/m3, which is significantly above the recommended remedial action level of 4 to 8 pCi/m3 if all leaks into the building are sealed (i.e., no more radon enters the building), how long will it take for the concentration of radon to reach the recommended level?
Answer #1

Answer-1)

Given:

The Radon 222 concentration in college classroom building = 1500 pCi/m3.

The recommended remedial action level of Radon 222 = [4-8 pCi/m3 ].

To calculate: The time it will take for the concentration of radon 222 to reach the recommended level.

Solve:

(i) To calculate The time it will take for the concentration of radon 222 to reach the recommended level in the building we will use the following radioactive law relation-

A1 = A0 x e(-λt).

Where A1 is the Activity (no. of radioactive deacy substance undergoes per second ) of radioactive substance after time t, A0 is the activity of the radioactive substance at time t = 0, and λ is the decay constant.

And the decay constant can be calculated using the following relation-

λ = ln(2) / t1/2 = 0.693 / t1/2.

Where t1/2 is the half-life of the radioactive substance.

Now We know that the halflife of radon 222 = 3.8 days so the decay constant for radon 222 gas is -

λ = 0.693 / t1/2.

λ = 0.693 / 3.8

λ = 0.1824 days-1.

(ii) So the initial activity (i.e at t= 0) of the radon 222 gas is 1500 pCi/m3, and let after time 't' the activity of the radon 222 gas is 8 pCi/m3 (taking upper limit of the recommended remedial action level of Radon 222 ), Then using the radioactive law relation-

A1 = A0 x e(-λt).

Substituting the values of λ , A1 and A0 in the above relation we get-

8 pCi/m3 = 1500 pCi/m3 x e(-0.1824 days-1 x t).

0.00533 = e(-0.1824 days-1 x t).

taking natural log both the side we get-

-5.23378 = -0.1824 days-1x t

t = 28.7 days

Similarly if we take the lower limit of the recommended remedial action level of Radon 222 (i.e 4 pCi/m3 ), we get-

4 pCi/m3 = 1500 pCi/m3 x e(-0.1824 days-1 x t).

0.00266 = e(-0.1824 days-1 x t).

taking natural log both the side we get-

-5.927 = -0.1824 days-1x t

t = 32.5 days

Answer: So The time it will take for the concentration of radon 222 to reach the recommended level is between 28.7 days to 32.5 days.

Hope you understood the answer.

Best Wishes!


answered by: ANURANJAN SARSAM